# Discrete Cosine Transform Based Decorrelated Attention for Vision Transformers

### v3

@@ Abstract
Central to the Transformer architectures’ effectiveness is the self-attention mechanism, a function that maps queries, keys, and values into a high-dimensional vector space. However, training the attention weights of queries, keys, and values is non-trivial from a state of random initialization. In this paper, we propose two methods. (i) We first address the initialization problem of Vision Transformers by introducing a simple, yet highly innovative, initialization approach utilizing discrete cosine transform (DCT) coefficients. Our proposed DCT-based attention initialization marks a significant gain compared to traditional initialization strategies; offering a robust foundation for the attention mechanism. Our experiments reveal that the DCT-based initialization enhances the accuracy of Vision Transformers in classification tasks. (ii) We also recognize that since DCT effectively decorrelates image information in the frequency domain, this decorrelation is useful for compression because it allows the quantization step to discard many of the higher-frequency components. Based on this observation, we propose a novel DCT-based compression technique for the attention function of Vision Transformers. Since high-frequency DCT coefficients usually correspond to noise, we truncate the high-frequency DCT components of the input patches. Our DCT-based compression reduces the size of weight matrices for queries, keys, and values. While maintaining the same level of accuracy, our DCT compressed Swin Transformers obtain a considerable decrease in the computational overhead.
Index Terms:
Discrete cosine transform, attention, vision transforms.
I Introduction
@@
In the last decade, convolutional neural networks (CNNs) have dominated image-processing tasks. However, inspired by the success of transformer architectures in natural language processing, Vision Transformers (ViTs) have redefined the norms of visual data interpretation in recent years. Self-attention mechanism, a key component of transformers, is pivotal in reshaping how models perceive and analyze visual information [1]. Unlike CNNs, ViTs view an image as a sequence of informative patches, adopting a more holistic approach. By treating image patches like “visual words”, ViTs can leverage powerful transformer architectures, originally designed for natural language processing, to learn global context and long-range dependencies. This novel approach has led to remarkable performance gains on a wide range of computer vision tasks, demonstrating the potential of ViTs to revolutionize the field.
@@
The initialization of self-attention weights is a critical step significantly impacting the training dynamics and performance of the Transformers. The core of Transformers is the self-attention mechanism, which includes three weight matrices: queries ($\mathbf{W}_{Q}$), keys ($\mathbf{W}_{K}$), and values ($\mathbf{W}_{V}$). Proper initialization of these matrices ensures that the model starts from a stable state, facilitating effective learning. The most common approach is to initialize the weights randomly, often using distributions like the normal or uniform distribution. For example, Xavier initialization [2] and He initialization [3] are popular choices. They are designed to keep the scale of the gradients roughly the same in all layers. Although less common, in some cases, biases or certain parts of the weight matrices are initialized to zeros [4, 5]. This is less common for queries, keys, and values matrices, as it can lead to symmetry-breaking issues during training [6].
@@
Efficiency is an ongoing concern in Transformers. While conventional initialization strategies are often suitable for generic applications, new algorithms are increasingly tailored to the specific characteristics and needs of Transformers. These advancements in initialization techniques contribute to the ongoing improvements in the performance and efficiency of Transformer-based models. Orthogonal initialization [7, 8], sparse initialization [9], variance scaling [10], pre-trained initialization [11, 12, 13], layer-wise, domain-specific and attention-specific initialization [14] are some of the new trends in initialization approaches. In this study, we propose a new family of algorithms for initialization strategy in the intersection of attention and domain-specific approaches: discrete cosine attention.
@@
Can DCT in Vision Transformer be a game changer? Discrete cosine transform (DCT) has been widely used in signal processing and data compression applications. It reduces the dimensionality of the input data while preserving its essential features. Transforming the input data into the frequency domain helps capture the most important features of the input data. In this paper, we propose two methods that apply DCT to the attention for vision transformers. First, we initialize the weight matrices for the queries, keys, and values computation as the DCT matrix. Our DCT-based initialization covers the entire spectrum utilizing the DCT basis vectors, and each weight vector has a different initial bandwidth. As a comparison, to cover the entire spectrum in the classical initialization, all the weight matrices are initialized as random numbers due to their independent generation. These random numbers can be treated as white noise. All the weight initialization starts with the same spectra. Second, we introduce a DCT-based compressed attention for vision transformers. With this compression methodology, compared to the vanilla Swin Transformers, the number of parameters and computational overhead of the Swin Transformers are reduced. Meanwhile, the accuracy of the models is comparable.
II Related works
II-A Vision Transformers
@@
Transformer architectures have merged as a dominant standard for natural language processing (NLP) tasks [1, 15]. Vision Transformer (ViT) [16] built a bridge between language and vision. It shows that pure transformers can also be applied directly to sequences of image patches for classification tasks. In other words, it partitions images into patches and treats them as word sequences. Then, a shifted windowing scheme was proposed in Swin Transformer [17] to obtain greater efficiency. This limits self-attention computation to non-overlapping local windows while allowing for cross-window connection. Later, to overcome the limitations of the heavyweights in ViTs, MobileViT [18] combined the strengths of MobileNets [19] and ViTs. It is a lightweight and low-latency network for mobile vision tasks. Moreover, gSwin [20] combined MLP-based architecture with Swin Transformer and attains superior accuracy with a smaller model size. Vision transformers are now supreme in various computer vision tasks such as super resolution [21, 22], compressive sensing [23, 24], image segmentation [25, 26, 27], and human mesh reconstruction [28].
II-B Frequency-based Neural Networks
@@
Orthogonal transforms such as Fourier transform [29, 30], Hadamard transform [31, 32], and discrete cosine Stockwell transform [33] have been used to improve the performance of the neural networks. In the orthogonal transform family, DCT is a mathematical technique commonly used in signal processing and image and video compression [34]. Compared to the Fourier transform, which is the standard frequency representation, DCT does not generate complex-valued numbers. Therefore, it is computationally more efficient. DCT-Former [35], for instance, applied DCT compression on the input sequence along the length channel before feeding into the self-attention layers to reduce the computational overhead in the NLP problems. Furthermore, inspired by JPEG compression, DCFormer [36] highlighted the visually significant signals by leveraging the input information compression on its frequency domain representation. These two DCT-based works and our DCT-based compression all aim to reduce the computational overhead. Therefore, we mainly compare our DCT-based compression method with them. DCT-Former [35] is designed for the NLP transformers and takes DCT along the width and height of the input tensor patches, while our method is designed for the vision transformers and it compresses the tensor along the channel axis. DCFormer [36] averages DCT coefficients via average pooling, which does not take advantage of the property that DCT components in low-frequency contain more important information. On the contrary, our DCT-based compression method retains the low-frequency components very accurately and discards the unnecessary extremely high-frequency components corresponding to noise. Under the same input image size, compared to the vanilla transformer models, DCFormer reduces the FLOPs and retains the number of parameters while sacrificing accuracy. Our DCT-based compression method, on the other hand, saves both trainable parameters and FLOPs and obtains on-par or better performance in addition to other benefits that our DCT-based compression method has. Furthermore, we introduce a DCT-based initialization and we have not found any related initialization in the literature.
II-C Initialization Methods for Self-Attention
@@
DS-Init [10] improves convergence in deep Transformers for NLP by reducing parameter variance during initialization and minimizing output variance of residual connections, facilitating better gradient flow. It pairs with a merged attention sublayer (MAtt) that combines average-based self-attention with encoder-decoder attention to reduce the computational cost. Additionally, Mimetic initialization [14] is a compute-free method that initializes the weights of self-attention layers by simulating the patterns found in pre-trained weights, resulting in improved Transformer performance. Moreover, weight selection [13] enables the initialization of smaller models by selecting a subset of weights from a larger pre-trained model, facilitating efficient knowledge transfer and faster training. This method enhances small model performance, reduces training time, and integrates with knowledge distillation, making it particularly valuable for resource-constrained environments. Furthermore, structured initialization [37] tackles the challenge of training Vision Transformers (ViTs) on small-scale datasets by introducing a structured initialization strategy inspired by the architectural biases of Convolutional Neural Networks (CNNs). While these methods focus on mimicking pre-trained patterns, selecting weights from larger models, or using CNN-inspired biases, our DCT-based initialization utilizes well-established DCT matrices to initialize attention weights.
II-D Efficient Multi-head Self-Attention
To reduce the memory and computational costs of dot-product attention, an efficient mechanism [38] in linear complexity was proposed. Besides, Hydra Attention [39] increased the number of heads to address the issue of the high computational complexity of the self-attention mechanism on large-sized images. Further, MetaFormer [40] verified that the overall architecture of Transformers, rather than the specific token mixer module, plays a more crucial role in the performance. Existing methods reduce self-attention complexity through linear attention, more heads, or architectural optimizations but ignore frequency-domain transformations. We propose DCT-based methods to enhance efficiency and expressiveness in ViTs.
III Methodology
In this section, we first review the multi-head self-attention (MSA) [17]. Then, we introduce our two proposed strategies to utilize the discrete cosine transform (DCT) in the attention function of the Swin Transformers. Finally, we discuss the motivation of our methods.
III-A Background: Multi-Head Self-Attention
@@
The overall structure of the vanilla MSA is drawn in Fig. 1a. To handle a 2D image tensor $\mathbf{I}\in\mathbb{R}^{H\times\ W\times C}$, where $H\times W$ are the resolution of the image and $C$ is the number of channels, we first split it into patches $\mathbf{X}\in\mathbb{R}^{N\times M^{2}\times C}$, where $M\times M$ is the resolution of each image patch and $N=\frac{WH}{M^{2}}$ is the resulting number of patches. After that, three linear layers ($\mathcal{L}_{Q}$, $\mathcal{L}_{K}$, $\mathcal{L}_{V}$) are applied on $\mathbf{X}$ to compute the queries, keys, and values $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times M^{2}\times C}$:
@@
where $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{C\times C}$ and $\mathbf{b}_{Q},\mathbf{b}_{K},\mathbf{b}_{V}\in\mathbb{R}^{1\times C}$ are trainable weights and biases. Next, to perform multi-head attention with $P$ heads, $\mathbf{Q,K,V}$ are split into $P$ slices as $\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}\in\mathbb{R}^{N\times M^{2}\times d}$ for $i=0,\ldots,P-1$, where $d=\frac{C}{P}$. The scaled dot-product attention $\mathbf{A}_{i}\in\mathbb{R}^{N\times M^{2}\times d}$ is computed as:
@@
where $\sigma(\cdot)$ is the Softmax function. $\mathbf{B}_{i}\in\mathbb{R}^{M^{2}\times M^{2}}$ is the relative position bias. Since the relative position along each axis lies in $[-M+1,M-1]$, values in $B_{i}$ are taken from a parameterized smaller-sized bias matrix $\hat{\mathbf{B}}_{i}\in\mathbb{R}^{(2M-1)\times(2M-1)}$. After that, $\{\mathbf{A}_{i}\}$ are concatenated to $\mathbf{A}\in\mathbb{R}^{N\times M^{2}\times C}$ to feed into a linear layer $\mathcal{L}_{O}$ to obtain the attention output $\mathbf{Y}\in\mathbb{R}^{N\times M^{2}\times C}$:
@@
(a) Vanilla / DCT-initialized MSA.
(b) DCT-Compressed MSA.
Figure 1: Multi-head Self-Attention. In (b), to reduce the computational overhead, IDCT and the final linear layer $\mathcal{L}_{O}$ are implemented as a single linear layer as no non-linearity exists in between.
III-B DCT-based Initialization for Attention
In the attention layer, the first three linear layers ($\mathcal{L}_{Q}$, $\mathcal{L}_{K}$, and $\mathcal{L}_{V}$) extract different features of the input patches $\mathbf{X}$. However, their weights $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ are trained simultaneously. Such training from scratch is difficult, especially when the dataset is insufficient and small-scale. On the other hand, DCT is a well-designed feature extractor. In our proposed DCT initialization strategy, one of $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, $\mathbf{W}_{V}\in\mathbb{R}^{C\times C}$ is initialized using the orthogonal type-II DCT matrix $\mathcal{D}$ [41]:
$\begin{split}&\{\mathbf{X}\mathcal{D}^{T}\}[k]\\ &=\begin{cases}\frac{\sqrt{2}}{C}\sum_{n=0}^{C-1}\mathbf{X}[n],&k=0,\\ \frac{2}{C}\sum_{n=0}^{C-1}\mathbf{X}[n]\cos\frac{(2n+1)k\pi}{2nC},&k=1,\ldots,C-1.\end{cases}\end{split}$ (4)
$\mathcal{D}$ can be obtained from applying DCT on an identity matrix $\mathbf{I}_{C\times C}$ because DCT is a linear transform. The remaining weight matrices and other trainable parameters are still initialized regularly. We don’t initialize multiple weight matrices to avoid the weight symmetry problem [6] caused by initializing the multiple weights as the same values, and our ablation study has verified this point.

### v5

# Discrete Cosine Transform-Based Decorrelated Attention for Vision Transformers

@@ Abstract
Self-attention is central to the success of Transformer architectures; however, learning the query, key, and value projections from random initialization remains challenging and computationally expensive. In this paper, we propose two complementary methods that leverage the Discrete Cosine Transform (DCT) to enhance the efficiency and performance of Vision Transformers. First, we address the initialization problem by introducing a simple yet effective DCT-based initialization strategy for self-attention, where projection weights are initialized using DCT coefficients. This structure-preserving approach consistently improves classification accuracy on the CIFAR-10 and ImageNet-1K benchmarks. Second, we propose a DCT-based attention compression technique that exploits the decorrelation properties of the frequency domain. By observing that high-frequency DCT coefficients typically correspond to noise, we truncate high-frequency components of the input patches, thereby reducing the dimensionality of the query, key, and value projections without sacrificing accuracy. Experiments on Swin Transformer models demonstrate that the proposed compression method achieves a substantial reduction in computational overhead while maintaining comparable performance. Code: https://github.com/NUBagciLab/DCT-Transformer.
1 Introduction
@@
In the last decade, convolutional neural networks (CNNs) have dominated image-processing tasks. However, inspired by the success of transformer architectures in natural language processing, Vision Transformers (ViTs) have redefined the norms of visual data interpretation in recent years. Self-attention mechanism, a key component of transformers, is pivotal in reshaping how models perceive and analyze visual information Vaswani et al. (2017).
Unlike CNNs, ViTs view an image as a sequence of informative patches, adopting a more holistic approach. By treating image patches like “visual words”, ViTs can leverage powerful transformer architectures, originally designed for natural language processing, to learn global context and long-range dependencies. This novel approach has led to remarkable performance gains on a wide range of computer vision tasks, demonstrating the potential of ViTs to revolutionize the field.
@@
The initialization of self-attention weights is a critical step significantly impacting the training dynamics and performance of the Transformers. The core of Transformers is the self-attention mechanism, which includes three weight matrices: queries ($\mathbf{W}_{Q}$), keys ($\mathbf{W}_{K}$), and values ($\mathbf{W}_{V}$). Proper initialization of these matrices ensures that the model starts from a stable state, facilitating effective learning. The most common approach is to initialize the weights randomly, often using distributions like the normal or uniform distribution. For example, Xavier initialization Glorot and Bengio (2010) and He initialization He et al. (2015) are popular choices. They are designed to keep the scale of the gradients roughly the same in all layers. Although less common, in some cases, biases or certain parts of the weight matrices are initialized to zeros Masood and Chandra (2012); Zhao et al. (2022). This is less common for queries, keys, and values matrices, as it can lead to symmetry-breaking issues during training Hu et al. (2019a).
@@
Efficiency is an ongoing concern in Transformers. While conventional initialization strategies are often suitable for generic applications, new algorithms are increasingly tailored to the specific characteristics and needs of Transformers. These advancements in initialization techniques contribute to the ongoing improvements in the performance and efficiency of Transformer-based models. Orthogonal initialization Saxe et al. (2013); Hu et al. (2019b), sparse initialization Martens and others (2010), variance scaling Zhang et al. (2019), pre-trained initialization Weiss et al. (2016); Bozinovski (2020); Xu et al. (2023), layer-wise, domain-specific and attention-specific initialization Trockman and Kolter (2023) are some of the new trends in initialization approaches. In this study, we propose a new family of algorithms for initialization strategy in the intersection of attention and domain-specific approaches: Discrete Cosine Transform-Based Decorrelated Attention.
@@
Can DCT in Vision Transformer be a game changer? DCT has been widely used in signal processing and data compression applications. It reduces the dimensionality of the input data while preserving its essential features. Transforming the input data into the frequency domain helps capture the most important features of the input data. In this paper, we propose two methods that apply DCT to the attention for vision transformers. First, we initialize the weight matrices for the queries, keys, and values computation as the DCT matrix. Our DCT-based initialization covers the entire spectrum utilizing the DCT basis vectors, and each weight vector has a different initial bandwidth. As a comparison, to cover the entire spectrum in the classical initialization, all the weight matrices are initialized as random numbers due to their independent generation. These random numbers can be treated as white noise. All the weight initialization starts with the same spectra. Second, we introduce a DCT-based compressed attention for vision transformers. With our compression methodology, compared to the vanilla Swin Transformers, the number of parameters and computational overhead of the Swin Transformers are reduced. Meanwhile, the accuracy of the models is comparable.
2 Related works
Vision Transformer
@@ 2 Related works
Transformer architectures have merged as a dominant standard for natural language processing tasks Vaswani et al. (2017). Vision Transformer (ViT) Dosovitskiy et al. (2020) built a bridge between language and vision. It shows that pure transformers can also be applied directly to sequences of image patches for classification tasks. In other words, it partitions images into patches and treats them as word sequences. Then, a shifted windowing scheme was proposed in Swin Transformer Liu et al. (2021) to obtain greater efficiency. This limits self-attention computation to non-overlapping local windows while allowing for cross-window connection. Later, to overcome the limitations of the heavyweights in ViTs, MobileViT Mehta and Rastegari (2021) combined the strengths of MobileNets Sandler et al. (2018) and ViTs. It is a lightweight and low-latency network for mobile vision tasks. Moreover, gSwin Go and Tachibana (2023) combined MLP-based architecture with Swin Transformer and attains superior accuracy with a smaller model size.
Frequency-based Neural Network
@@
Orthogonal transforms such as the Fourier transform Chi et al. (2020); Buchholz and Jug (2022); Patro et al. (2025) and Hadamard transform Pan et al. (2023) have been used to improve the performance of the neural networks. In the orthogonal transform family, DCT is a mathematical technique commonly used in signal processing and image and video compression. Compared to the Fourier transform, which is the standard frequency representation, DCT does not generate complex-valued numbers. Therefore, it is computationally more efficient. DCT-Former Scribano et al. (2023), for instance, applied DCT compression on the input sequence along the length channel before feeding it into the self-attention layers to reduce the computational overhead in the NLP problems. Furthermore, inspired by JPEG compression, DCFormer Li et al. (2023) highlighted the visually significant signals by leveraging the input information compression on its frequency domain representation. These two DCT-based works and our DCT-based compression all aim to reduce the computational overhead. Therefore, we mainly compare our DCT-based compression method with them. DCT-Former Scribano et al. (2023) is designed for the NLP transformers and takes DCT along the width and height of the input tensor patches, while our method is designed for the vision transformers and it compresses the tensor along the channel axis. DCFormer Li et al. (2023) averages DCT coefficients via average pooling, which does not take advantage of the property that DCT components in low-frequency bands contain more important information. On the contrary, our DCT-based compression method retains the low-frequency components very accurately and discards the unnecessary extremely high-frequency components corresponding to noise. Under the same input image size, compared to the vanilla transformer models, DCFormer reduces the FLOPs and retains the number of parameters while sacrificing accuracy. Our DCT-based compression method, on the other hand, saves both trainable parameters and FLOPs and obtains on-par or better performance in addition to other benefits that our DCT-based compression method has. Furthermore, we introduce a DCT-based initialization and we have not found any related initialization in the literature.
Self-Attention Initialization
@@
DS-Init Zhang et al. (2019) improves convergence in deep Transformers for NLP by reducing parameter variance during initialization and minimizing output variance of residual connections, facilitating better gradient flow. It pairs with a merged attention sublayer (MAtt) that combines average-based self-attention with encoder-decoder attention to reduce the computational cost. Additionally, Mimetic initialization Trockman and Kolter (2023) is a compute-free method that initializes the weights of self-attention layers by simulating the patterns found in pre-trained weights, resulting in improved Transformer performance. Moreover, weight selection Xu et al. (2023) enables the initialization of smaller models by selecting a subset of weights from a larger pre-trained model, facilitating efficient knowledge transfer and faster training. This method enhances small model performance, reduces training time, and integrates with knowledge distillation, making it particularly valuable for resource-constrained environments. Furthermore, structured initialization Zheng et al. (2024) tackles the challenge of training vision transformers on small-scale datasets by introducing a structured initialization strategy inspired by the architectural biases of Convolutional Neural Networks (CNNs). While these methods focus on mimicking pre-trained patterns, selecting weights from larger models, or using CNN-inspired biases, our DCT-based initialization utilizes well-established DCT matrices to initialize attention weights.
Efficient Self-Attention
A variety of approaches have been proposed to reduce the memory and computational costs of dot-product attention, including linear-complexity attention mechanisms Shen et al. (2021), increasing attention heads as in Hydra Attention Bolya et al. (2022), and architectural rethinking exemplified by MetaFormer Yu et al. (2022), which shows that overall architecture often matters more than the specific token mixer. MicroViT Setyawan et al. (2025) further targets resource-constrained settings by introducing an Efficient Single Head Attention mechanism that reduces self-attention cost through group convolution and partial channel processing. Despite these advances, existing methods mainly pursue efficiency through linearization, head scaling, or architectural design, while largely ignoring frequency-domain transformations. We propose DCT-based methods to enhance both efficiency and expressiveness in vision transformers.
3 Methodology
3.1 Background: Multi-Head Self-Attention
@@
To handle a 2D image tensor $\mathbf{I}\in\mathbb{R}^{H\times\ W\times C}$, the multi-head self-attention (MSA) module first splits it into patches $\mathbf{X}\in\mathbb{R}^{N\times M^{2}\times C}$, where $H\times W$ are the resolution of the image and $C$ is the number of channels, $M\times M$ is the resolution of each image patch and $N=\frac{WH}{M^{2}}$ is the resulting number of patches. After that, three linear layers ($\mathcal{L}_{Q}$, $\mathcal{L}_{K}$, $\mathcal{L}_{V}$) are applied on $\mathbf{X}$ to compute the queries, keys, and values $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times M^{2}\times C}$:
@@
where $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{C\times C}$ and $\mathbf{b}_{Q},\mathbf{b}_{K},\mathbf{b}_{V}\in\mathbb{R}^{1\times C}$ are trainable weights and biases. Then, to perform $P$-head MSA, $\mathbf{Q,K,V}$ are split into $P$ slices as $\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}\in\mathbb{R}^{N\times M^{2}\times d}$ for $i=0,\ldots,P-1$, $d=\frac{C}{P}$. The scaled dot-product attention $\mathbf{A}_{i}\in\mathbb{R}^{N\times M^{2}\times d}$ is computed as:
@@
where $\sigma(\cdot)$ is the Softmax function. $\mathbf{B}_{i}\in\mathbb{R}^{M^{2}\times M^{2}}$ is the relative position bias. Since the relative position along each axis lies in [M+1,M1][-M+1,M-1], values in $B_{i}$ are taken from a parameterized smaller-sized bias matrix $\hat{\mathbf{B}}_{i}\in\mathbb{R}^{(2M-1)\times(2M-1)}$. After that, $\{\mathbf{A}_{i}\}$ are concatenated to $\mathbf{A}\in\mathbb{R}^{N\times M^{2}\times C}$ to feed into a linear layer $\mathcal{L}_{O}$ to obtain the attention output $\mathbf{Y}\in\mathbb{R}^{N\times M^{2}\times C}$:
@@
(a) DCT-Initializion.
(b) DCT-Compression.
Figure 1: DCT-based initialization and DCT-based compressed MSA. In (a), the DCT-Initialized MSA uses the same structure as the vanilla MSA, with weights initialized as the DCT matrix. In (b), to reduce the computational overhead, IDCT and the final linear layer $\mathcal{L}_{O}$ are implemented as a single linear layer because there exists no non-linearity in between.
3.2 DCT-based Initialization for Attention
In the attention layer, the first three linear layers ($\mathcal{L}_{Q}$, $\mathcal{L}_{K}$, and $\mathcal{L}_{V}$) extract different features of the input patches $\mathbf{X}$. However, their weights $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ are trained simultaneously. Such training from scratch is difficult, especially when the dataset is insufficient and small-scale. On the other hand, DCT is a well-designed feature extractor. In our proposed DCT initialization strategy, one of $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, $\mathbf{W}_{V}\in\mathbb{R}^{C\times C}$ is initialized using the orthogonal type-II DCT matrix $\mathcal{D}$ Ahmed et al. (1974):
[…]
$\mathcal{D}$ can be obtained from applying DCT on an identity matrix $\mathbf{I}_{C\times C}$ because DCT is a linear transform. The remaining weight matrices and other trainable parameters are still initialized regularly. We don’t initialize multiple weight matrices to avoid the weight symmetry problem Hu et al. (2019a) caused by initializing the multiple weights as the same values, and our ablation study has verified this point.